If 3 melons sell for $0.42, how much would 7 melons cost?
Answer by TimothyLamb(4379)

0.42/3 = x/7
x = (0.42/3)*7
x = $0.98
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php